OpenStudy (anonymous):

find the difference quotient f(x+h)-f(x)/h for the function f (x)=-3x^2+7x-2. a) -6x-6h+7 b) -6x-3h+7 c) -6x-3h+7-6x^2+14x-4/hx d) -6x-6h+5 e) none of the above

OpenStudy (zarkon):

b

OpenStudy (anonymous):

explain please?

OpenStudy (zarkon):

\[\frac{f(x+h)-f(x)}{h}=\frac{-3(x+h)^2+7(x+h)-2-(-3x^2+7x-2)}{h}\] \[=\frac{-3(x+h)^2+7(x+h)-2+3x^2-7x+2}{h}\] \[=\frac{-3x^2-6xh-3h^2+7x+7h-2+3x^2-7x+2}{h}\] \[=\frac{-6xh-3h^2+7h}{h}=-6x-3h+7\]